OpenStudy (anonymous):

Write 2 - 7i in polar coordinates on the complex plane

OpenStudy (anonymous):

TJ , you are here, help him, please

OpenStudy (anonymous):

Write 2 - 7i in polar coordinates on the complex plane.

terenzreignz (terenzreignz):

I came here to watch you, @Hoa :/ Anyway, @Escobar12 to get the polar coordinates of a complex number in rectangular (a + bi) form, you need two numbers, \(r\) and \(\theta\) Use these equations \[\huge r = \sqrt{a^2+b^2}\]\[\huge \tan (\theta) = \frac{b}a\]

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

@Escobar12 waaaha, you are tooooo fast . I didn't get what he mean yet @terenzreignz you know that I cannot work under that pressure, teasing me???

terenzreignz (terenzreignz):

I'm not yet quite finished, @Escobar12 Once you get these two numbers, write like this... \[\huge a+bi = r[\cos(\theta) +i\sin(\theta)]\]

OpenStudy (anonymous):

continue TJ , it's not easy, give him more explanation, please.

terenzreignz (terenzreignz):

Okay, an example... \[\huge 1 + i\sqrt3\] To get r, we use the formula \[\huge r = \sqrt{1^2+(\sqrt3)^2}=\sqrt{1+3}=\sqrt4=2\]\[\huge \tan(\theta) =\frac{\sqrt3}{1}=\sqrt3\\\ \huge \theta = 60^o \left(or \quad \theta = \frac{\pi} {3}\right)\] so finally, we get. \[\huge 1+i\sqrt3=2\left[\cos\left(\frac{\pi}{3}\right)+i\sin\left(\frac{\pi}{3}\right)\right]\]

OpenStudy (anonymous):

ooohkay so what if i use the numbers to Find (1 + i)8 using DeMoivre's Theorem

terenzreignz (terenzreignz):

First, put 1+i in polar form.

OpenStudy (anonymous):

and i use the first equation you gave me right?

OpenStudy (anonymous):

@terenzreignz knock, knock @Escobar12 do your stuff, no need to wait for confirmation. he will check yours then

terenzreignz (terenzreignz):

I'm here. And yes, @Escobar12 Go ahead and use the equations I gave you :)

OpenStudy (anonymous):

@terenzreignz let me take it over, ok, and you must go, I will call you back to check, ok?

OpenStudy (anonymous):

@ Escoba12, tell me what part you cannot see the link?

terenzreignz (terenzreignz):

Aching for some adventure, @Hoa ? Okay then :D

OpenStudy (anonymous):

\[\huge a+bi = r[\cos(\theta) +i\sin(\theta)\]\It's the formula you need to transpose from complex to polar. yours is 2 -7i, it means your a = 2, your b = -7 got it? and from right hand side, the easier way to get is = rcos (theta) + r *i *sin(theta)

OpenStudy (anonymous):

now, balance both side, real part from LHS = real part from RHS. what does it mean? It means 2 = r cos theta.

OpenStudy (anonymous):

got it? the imaginary part is the same, you have -7i = r i sin(theta) that means -7 = r sin (theta) you have 2 new equations, 2 variables r and theta, solve for them to get r, and theta. then put them into the polar plane.

OpenStudy (anonymous):

ha!!! I'm out of battery.!!!

OpenStudy (anonymous):

thanks for being very helpful(:

OpenStudy (anonymous):

@terenzreignz back here

OpenStudy (anonymous):

@Mertsj please, rescue.

OpenStudy (anonymous):

i have to go now, let other help

OpenStudy (mertsj):

Looks good to me. In summary 1. Draw a picture  2. Use the Pythagorean Theorem to find the hypotenuse (polar radius) 3. Use the tangent function to find the angle theta 4. Put the newly discovered information into the polar form of a complex number which is \[r(\cos \theta +i \theta \sin \theta)\]

terenzreignz (terenzreignz):

Sorry for my late arrival... Lag...

OpenStudy (anonymous):

so the answer in polar form is \[\sqrt{i}\]=(cos(pi/2+i sin(pi/2))^1/2

OpenStudy (mertsj):